# mOSCAR: A Large-scale Multilingual and Multimodal Document-level Corpus

### v1

@@ Abstract
Multimodal Large Language Models (mLLMs) are trained on a large amount of text-image data. While most mLLMs are trained on caption-like data only, Alayrac et al. (2022) showed that additionally training them on interleaved sequences of text and images can lead to the emergence of in-context learning capabilities. However, the dataset they used, M3W, is not public and is only in English. There have been attempts to reproduce their results but the released datasets are English-only. In contrast, current multilingual and multimodal datasets are either composed of caption-like only or medium-scale or fully private data. This limits mLLM research for the 7,000 other languages spoken in the world. We therefore introduce mOSCAR, to the best of our knowledge the first large-scale multilingual and multimodal document corpus crawled from the web. It covers 163 languages, 315M documents, 214B tokens and 1.2B images. We carefully conduct a set of filtering and evaluation steps to make sure mOSCAR is sufficiently safe, diverse and of good quality. We additionally train two types of multilingual model to prove the benefits of mOSCAR: (1) a model trained on a subset of mOSCAR and captioning data and (2) a model train on captioning data only. The model additionally trained on mOSCAR shows a strong boost in few-shot learning performance across various multilingual image-text tasks and benchmarks, confirming previous findings for English-only mLLMs. The dataset can be accessed here. [^1] https://oscar-project.github.io/documentation/versions/mOSCAR/
@@ 1 Introduction
This motivated us to collect and release the first large-scale multilingual and multimodal document dataset derived from Common Crawl. [^2] https://commoncrawl.org/. The Common Crawl Foundation is a non-profit organization that crawls the web on a monthly basis. Our dataset, multimodal OSCAR (mOSCAR), follows the OSCAR initiative (Ortiz Suárez et al., 2019; Abadji et al., 2021, 2022) and covers 315M documents in 163 languages, 214B tokens and 1.2B images. Figure 1 shows an example of a document, more can be found in Appendix A.3. We carry out extensive filtering to increase its safety and quality. To prove mOSCAR’s utility, we train a multilingual OpenFlamingo (Awadalla et al., 2023) from a Gemma-2B language model (Gemma Team et al., 2024) on a subset of mOSCAR and captioning data from LAION-400M (Schuhmann et al., 2021), recaptioned with BLIP (Li et al., 2022), filtered with CLIP (Radford et al., 2021) and translated with NLLB (Costa-jussà et al., 2022). We compare against a similar model trained on captioning data only and show we obtain a strong boost in few-shot learning, confirming previous findings for English (Alayrac et al., 2022; McKinzie et al., 2024; Laurençon et al., 2024). The dataset and models will be made publicly available.
@@ 3.3 Text-only filtering
We conduct several types of per-language deduplication at different levels, as this has been shown to improve training efficiency (Abbas et al., 2023). First, we keep unique documents only by removing exact duplicates at the document level. We also remove exact duplicates of text nodes within the same document (4% of text nodes) and near-duplicate text nodes (1% of text nodes) by computing the Levenshtein ratio (Levenshtein, 1966) between all text nodes within the same document and applying a threshold of 0.95. If near-duplicates are found, we keep the first one in the document. Finally, we conduct per language near-deduplication at the document level with MinHashLSH (Broder, 1997; Gionis et al., 1999) following Smith et al. (2022), removing on average 19% of documents: [^5] With some disparity among languages as we found more duplicates for low- than high-resource languages. we turn documents into hashing vectors, compute min hashes from these vectors and perform Locality Sensitive Hashing to remove duplicates [^6] We performed this using the datasketch python library. (see Appendix A.5 for more details).
@@ 3.4 Image-only filtering
We use multiple NSFW automatic models to remove as much unsafe content as possible. We first combine two NSFW detectors: nsfw-detector (Laborde, ), a 5-class classifier with a MobileNet (Howard et al., 2017) backbone fine-tuned on 60GB of annotated data and NudeNet, [^7] https://github.com/vladmandic/nudenet an object detector trained to detect different types of nudity in images. We combined the two models as we found the first to be gender-biased while the second gives a large number of false positives for non-human images. Concretely, we consider an image an NSFW candidate if the sum of the probabilities for the classes ‘porn’ and ‘hentai’ is superior to 0.8 using nsfw-detector. We then tag the image as NSFW if one of the sensitive ‘exposed’ classes of NudeNet gets a probability superior to 0.5. We additionally use Safer by Thorn [^8] https://safer.io/, a private pornography detector, and tag the image as NSFW if the probability of the class ‘pornography’ is superior to 0.8. If a document contains an image with an NSFW tag, we remove the entire document from the dataset, which removes 0.5% of images. We manually inspecting 1,000 images of the remaining data and found no NSFW content. We manually inspected 1,000 images of the removed content and found 63.4% of NSFW images.
@@
Child Sexual Abuse Material (CSAM) is widespread on the internet and is therefore likely to be found in such a large-scale dataset crawled from the web. Removing CSAM is challenging as there is no training data nor open-source detection models available as these could be used in a harmful way. We again rely on Safer, a proprietary 3-class classifier trained to detect CSAM and pornography content from images. We tag the image as CSAM if the probability of the class CSAM is superior to 0.4 to favour recall over precision. As mentioned above, if a document contains an image with a CSAM tag, we remove it from the dataset. This step removes 0.07% of the images.
@@ 3.6 Text-image joint filtering
Our aim is to obtain truly multimodal documents where all images are related to at least one of the text nodes in some way [^10] We do not limit ourselves to caption-like relation and instead allow all types of text-image relation. and vice versa. We choose to apply joint text-image filtering to discard images and/or text nodes that are irrelevant to the rest of the document (e.g. the case of ads and website functionalities). To do this, we use NLLB-SIGLIP111111siglip-base-patch16-224 as vision encoder and nllb-distilled-600M as text encoder. (Visheratin, 2023), a multilingual version of SIGLIP (Zhai et al., 2023) trained with the encoder of NLLB (Costa-jussà et al., 2022), which covers all mOSCAR languages. [^12] We use the open-clip (Ilharco et al., 2021) model version and the transformers (Wolf et al., 2020) library. We compute cosine similarity scores between all images and all paragraphs [^13] We refer to paragraph as the text content in a DOM text node. within a same document. To remove irrelevant text nodes or images in a document, we mimic a text-image retrieval task, which means we avoid using arbitrary cosine similarity thresholds for each language and can reduce length biases and those in favour of caption-like paragraphs. For each candidate pair we randomly sample 63 negative images and 63 negative similar-length paragraphs from the same language but other documents. We tag the text node (resp. image) as valid if the cosine similarity of the pair is among the top 8 of the text-to-image (resp. image-to-text) similarity scores computed with the candidate text node (resp. image) and all the negative images (resp. text nodes). This means that we tag the text node (resp. image) as valid if it has a significantly higher score than a score computed with a random image (resp. text) for at least one of the images (resp. text node) in the document. We then discard text nodes and images not tagged as valid (on average 35% of the DOM text nodes and 10% of the images within a document). After this filtering step, we apply additional text-only filters to keep documents superior to 100 bytes.
@@ 4 Multimodal Open Super-large Crawled Aggregated coRpus (mOSCAR)
mOSCAR is extracted from three Common Crawl dumps from 2023. Due to computational constraints and in order to extract a maximum number of documents for low-resource languages, we extracted all languages from the first dump only. We removed the 6 most high-resource languages from the second dump and only extracted the languages with fewer than 1M documents for the last dump. Table 1 shows a distribution of the total number of languages and their number of documents.
To avoid data poisoning (Carlini et al., 2024), we release a hash (sha512) with each mOSCAR image.
#documents 10M 5M 1M 500K 200K 50K 10K 5K 1K
#languages 10 15 38 49 58 82 129 142 163
@@
Figure 2: Distributions of numbers of tokens and images per document
mOSCAR is composed of 315M documents (214B tokens, 1.2B images) from 163 languages. Figure 2 shows the distribution of images and tokens per document and their joint distribution. As shown in Figure 2(a), the mean and median number of images per document is 2 and 3.80.
@@ 4.1.1 Text content
By contruction, mOSCAR is diverse in terms of number of languages, so we focus on the diversity of mOSCAR’s English documents and compare against mmc4 (Zhu et al., 2023), OBELICS (Laurençon et al., 2023) and the English subset of WIT (Srinivasan et al., 2021). We compute the Vendi score (Friedman and Dieng, 2023) on a set of SimCSE embeddings (Gao et al., 2021) with a RoBERTa encoder (Liu et al., 2019) to evaluate the content diversity. Since embedding-based diversity metrics target content diversity well but are less relevant for lexical diversity (Tevet and Berant, 2021), we measure lexical diversity via the distinct $n$-gram ratio (Li et al., 2016).
@@ 4.1.2 Image diversity
(a) Comparison of different datasets.
@@
(b) mOSCAR (English vs. any language).
Table 3: Average Vendi score ($\pm$ standard error) of images.
We compute the Vendi Score on random samples of images for different datasets, comparing the images from English mOSCAR documents with those from Conceptual Captions (Changpinyo et al., 2021), LAION-400M (Schuhmann et al., 2021) and WIT (Srinivasan et al., 2021). We represent each image by its SigLIP [^14] We use siglip-base-patch16-224. (Zhai et al., 2023) embedding and compute the Vendi score on batches of size 65,536 and a total of 1M images for each dataset. In Table 3(a), we notice that the set of images in mOSCAR documents are more diverse than images from WIT documents but less diverse than LAION-400M.
@@
We also compare the diversity of images from English documents and of images sampled from documents of any language (English included). We use multilingual SigLIP (Chen et al., 2023a) trained on WebLI (Chen et al., 2023b) to compute image embeddings used to get the Vendi score. We again use a batch of size 65,536 and a total of 3M images, and we do not sample multiple images from a same document. For the multilingual setting, we randomly sample 50 languages and an equal number of images for each language to build the batch. As we did not do any image deduplication across languages, we could expect to have less diversity in the multilingual setting. However, Table 3(b) shows that the set of images is on average more diverse when sampled from all documents than from English-only documents. This means that the distribution of images is not exactly the same across languages, potentially due to cultural differences.
5 Training a multilingual multimodal language model
We train a multilingual Flamingo-like model on mOSCAR. As adding captioning data to training data has been shown to improve zero-shot performance (McKinzie et al., 2024), we additionally train on LAION-400M, which we re-captioned using BLIP (Li et al., 2022), filtered with CLIP score (Radford et al., 2021) and translated using distilled NLLB-600M (Visheratin, 2023) following the proportion of languages found in mOSCAR. We use Gemma-2B (Gemma Team et al., 2024) as the underlying language model and we train the model on 35M mOSCAR documents and 70M randomly sampled image-text pairs. We also train a model on 300M image-text pairs as a comparison baseline. We additionally compare with OpenFlamingo-3B-MPT (Awadalla et al., 2023) as the translate-test baseline. The full list of languages for training and the implementation details can be found in Appendix A.5.
5.1 Evaluation setup
5.2 Results
Tables 4 and 5 show the average results across all languages. Full results are available in Appendix A.6. We notice that the multilingual OpenFlamingo trained additionally on mOSCAR gets better results than the model trained on captioning data only while having seen fewer image-text pairs during training. More importantly, when increasing the number of few-shot examples from 0 to 16, it sees gains of on average +8.19 points on VQA benchmarks and +16.07 CideR points on captioning benchmarks. In contrast, the model trained on text-image pairs only sees gains of +2.82 and +9.08 points respectively. In cross-modal machine translation, the model additionally trained on interleaved data is again far better than the one trained on just captioning data, which is not able to translate the Multi30k benchmark at all. [^15] Most of the time, the model is not able to follow the prompt and only outputs the end of sequence token. Moreover, mOSCAR helps the model to learn to zero-shot disambiguate translations as shown by the improved average score on CoMMuTE (63.75) compared to the model trained on captions only (61.36).
Multilingual Open Flamingo trained on mOSCAR & text-image pairs is also better than OpenFlamingo 3B MPT evaluated on translate test benchmarks. However, we obtain the best results (except for MaXM) by evaluating our multilingual Open Flamingo on the translate-test benchmarks since the underlying language model (Gemma-2B) is far better in English than other languages. We also notice that all models struggle with reasoning classification tasks (MaRVL, XVNLI) where they obtain scores close to random guessing.
@@ 5.2 Results
Table 4: Results averaged over all languages.

### v2

@@ Abstract
Multimodal Large Language Models (mLLMs) are trained on a large amount of text-image data. While most mLLMs are trained on caption-like data only, Alayrac et al. (2022) showed that additionally training them on interleaved sequences of text and images can lead to the emergence of in-context learning capabilities. However, the dataset they used, M3W, is not public and is only in English. There have been attempts to reproduce their results but the released datasets are English-only. In contrast, current multilingual and multimodal datasets are either composed of caption-like only or medium-scale or fully private data. This limits mLLM research for the 7,000 other languages spoken in the world. We therefore introduce mOSCAR, to the best of our knowledge the first large-scale multilingual and multimodal document corpus crawled from the web. It covers 163 languages, 303M documents, 200B tokens and 1.15B images. We carefully conduct a set of filtering and evaluation steps to make sure mOSCAR is sufficiently safe, diverse and of good quality. We additionally train two types of multilingual model to prove the benefits of mOSCAR: (1) a model trained on a subset of mOSCAR and captioning data and (2) a model trained on captioning data only. The model additionally trained on mOSCAR shows a strong boost in few-shot learning performance across various multilingual image-text tasks and benchmarks, confirming previous findings for English-only mLLMs. The dataset is released under the Creative Commons CC BY 4.0 license and can be accessed here. [^1] https://oscar-project.github.io/documentation/versions/mOSCAR/
mOSCAR: A Large-scale Multilingual
and Multimodal Document-level Corpus
Matthieu Futeralthanks: Correspondence to matthieu.futeral@inria.fr   1,2    Armel Zebaze1,4    Pedro Ortiz Suarez5    Julien Abadji1
Rémi Lacroix3, 6    Cordelia Schmid1,2    Rachel Bawden1    Benoît Sagot1 1Inria     2Département d’informatique de l’ENS, CNRS, PSL Research University 3Institut du développement et des ressources en informatique scientifique, CNRS 4Sorbonne Université, Paris, France     5Common Crawl Foundation     6Université Paris-Saclay
@@ 1 Introduction
This motivated us to collect and release the first large-scale multilingual and multimodal document dataset derived from Common Crawl. [^2] https://commoncrawl.org/. The Common Crawl Foundation is a non-profit organization that crawls the web on a monthly basis. Our dataset, multimodal OSCAR (mOSCAR), follows the OSCAR initiative (Ortiz Suárez et al., 2019; Abadji et al., 2021, 2022) and covers 303M documents in 163 languages, 200B tokens and 1.15B images. Figure 1 shows an example of a document, more can be found in Appendix A.3. We carry out extensive filtering to increase its safety and quality. To prove mOSCAR’s utility, we train a multilingual OpenFlamingo (Awadalla et al., 2023) from a Gemma-2B language model (Gemma Team et al., 2024) on a subset of mOSCAR and captioning data from LAION-400M (Schuhmann et al., 2021), recaptioned with BLIP (Li et al., 2022), filtered with CLIP (Radford et al., 2021) and translated with NLLB (Costa-jussà et al., 2022). We compare against a similar model trained on captioning data only and show we obtain a strong boost in few-shot learning, confirming previous findings for English (Alayrac et al., 2022; McKinzie et al., 2024; Laurençon et al., 2024). mOSCAR can be accessed here  [^3] https://huggingface.co/datasets/oscar-corpus/mOSCAR.
@@ 3.3 Text-only filtering
We conduct several types of per-language deduplication at different levels, as this has been shown to improve training efficiency (Abbas et al., 2023). First, we keep unique documents only by removing exact duplicates at the document level. We also remove exact duplicates of text nodes within the same document (4% of text nodes) and near-duplicate text nodes (1% of text nodes) by computing the Levenshtein ratio (Levenshtein, 1966) between all text nodes within the same document and applying a threshold of 0.95. If near-duplicates are found, we keep the first one in the document. Finally, we conduct per language near-deduplication at the document level with MinHashLSH (Broder, 1997; Gionis et al., 1999) following Smith et al. (2022), removing on average 19% of documents: [^6] With some disparity among languages as we found more duplicates for low- than high-resource languages. we turn documents into hashing vectors, compute min hashes from these vectors and perform Locality Sensitive Hashing to remove duplicates [^7] We performed this using the datasketch python library. (see Appendix A.6.1 for more details).
Toxicity filtering
Toxic content targeting individuals or groups of people is widespread on the internet and can therefore be found in large-scale web-crawled datasets like mOSCAR without appropriate filtering steps. To alleviate this issue, we apply the same method used by Costa-jussà et al. (2022) and remove documents from mOSCAR based on the presence of a list of “toxic” words for each language [^8] The list of these words for each language can be found here: https://github.com/facebookresearch/flores/tree/main/toxicity. As some words in the list can also be used in a non-toxic way based on the context (e.g.: ‘breast’ in English), we tag the document as toxic and remove it from mOSCAR if it contains at least two distinct words in the list. This filtering step removes 0.95% of the documents for very high-resource languages (>5M documents), 2.13% for high-resource languages (<5M, >500K), 0.47% for mid-resource languages (<500K, >50K) and 0.64% for low-resource languages (< 50K). When manually analysing 1,000 random documents removed by this filtering step in each of the 2 (high-resource) languages we are native speakers of (English and French), we found 568 documents with toxic content.
Personal Identifiable Information
Personal Identifiable Information (PII) can be found in large-scale web-crawled datasets, we therefore conducted an additional filtering step to replace all detected PII by place holder strings using regular expressions (Appendix A.6.2 for more details). Concretely, we replaced all detected email addresses, phone numbers, credit card numbers, IP addresses and passport numbers. Moreover, it was shown that CommonCrawl contains a non neglictible part of API keys in its content  [^9] https://trufflesecurity.com/blog/[…]. We therefore scanned the dataset with the trufflehog tool  [^10] https://github.com/trufflesecurity/trufflehog to track down residual API keys that could have passed previous filters. We found $\sim$200K positive matches and manually check a random sample of 1K positive matches. We found only 2 of them to potentially be API keys, other matches are mainly noisy text nodes not related to PII. We removed the 200K text nodes from mOSCAR.
@@ 3.4 Image-only filtering
We use multiple NSFW automatic models to remove as much unsafe content as possible. We first combine two NSFW detectors: nsfw-detector (Laborde, ), a 5-class classifier with a MobileNet (Howard et al., 2017) backbone fine-tuned on 60GB of annotated data and NudeNet, [^11] https://github.com/vladmandic/nudenet an object detector trained to detect different types of nudity in images. We combined the two models as we found the first to be gender-biased while the second gives a large number of false positives for non-human images. Concretely, we consider an image an NSFW candidate if the sum of the probabilities for the classes ‘porn’ and ‘hentai’ is superior to 0.8 using nsfw-detector. We then tag the image as NSFW if one of the sensitive ‘exposed’ classes of NudeNet gets a probability superior to 0.5.
If a document contains an image with an NSFW tag, we remove the entire document from the dataset, which removes 0.5% of images. We manually inspecting 1,000 images of the remaining data and found no NSFW content. We manually inspected 1,000 images of the removed content and found 63.4% of NSFW images.
@@
Child Sexual Abuse Material (CSAM) is widespread on the internet and is therefore likely to be found in such a large-scale dataset crawled from the web. Removing CSAM is challenging as there is no training data nor open-source detection models available as these could be used in a harmful way. We rely on Thorn’s CSAM classifier [^12] https://safer.io/, a proprietary classifier trained to detect CSAM content in images. We tag the image as CSAM if the probability of the class CSAM is superior to 0.4 to favour recall over precision. As mentioned above, if a document contains an image with a CSAM tag, we remove it from the dataset. This step removes 0.07% of the images.
@@
Personal Identification Information
To protect PII in images, we use a lightweight face detector [^14] https://github.com/Linzaer/Ultra-Light-Fast-Generic-Face-Detector-1MB and apply a threshold of 0.99 to detect faces in the images. We apply such a high threshold as we found the model to be biased towards detecting faces with high probability in images without any human. For each image in mOSCAR, we distribute the bounding boxes of the detected faces so that users can blur them when downloading the images. More details are provided in Appendix A.6.2.
@@ 3.6 Text-image joint filtering
Our aim is to obtain truly multimodal documents where all images are related to at least one of the text nodes in some way [^15] We do not limit ourselves to caption-like relation and instead allow all types of text-image relation. and vice versa. We choose to apply joint text-image filtering to discard images and/or text nodes that are irrelevant to the rest of the document (e.g. the case of ads and website functionalities).
To do this, we use NLLB-SIGLIP161616siglip-base-patch16-224 as vision encoder and nllb-distilled-600M as text encoder. (Visheratin, 2023), a multilingual version of SIGLIP (Zhai et al., 2023) trained with the encoder of NLLB (Costa-jussà et al., 2022), which covers all mOSCAR languages. [^17] We use the open-clip (Ilharco et al., 2021) model version and the transformers (Wolf et al., 2020) library. We compute cosine similarity scores between all images and all paragraphs [^18] We refer to paragraph as the text content in a DOM text node. within a same document. To remove irrelevant text nodes or images in a document, we mimic a text-image retrieval task, which means we avoid using arbitrary cosine similarity thresholds for each language and can reduce length biases and those in favour of caption-like paragraphs. For each candidate pair we randomly sample 63 negative images and 63 negative similar-length paragraphs from the same language but other documents. We tag the text node (resp. image) as valid if the cosine similarity of the pair is among the top 8 of the text-to-image (resp. image-to-text) similarity scores computed with the candidate text node (resp. image) and all the negative images (resp. text nodes). This means that we tag the text node (resp. image) as valid if it has a significantly higher score than a score computed with a random image (resp. text) for at least one of the images (resp. text node) in the document. We then discard text nodes and images not tagged as valid (on average 35% of the DOM text nodes and 10% of the images within a document).
After this filtering step, we apply additional text-only filters to keep documents superior to 100 bytes. We also reapply the open-lid language detector (Burchell et al., 2023) as described in Section 3.2 as we found the last filtering step to change the major language of some documents.
@@ 4 Multimodal Open Super-large Crawled Aggregated coRpus (mOSCAR)
mOSCAR is extracted from three Common Crawl dumps from 2023. Due to computational constraints and in order to extract a maximum number of documents for low-resource languages, we extracted all languages from the first dump only. We removed the 6 most high-resource languages from the second dump and only extracted the languages with fewer than 1M documents for the last dump. Table 1 shows a distribution of the total number of languages and their number of documents. To avoid data poisoning (Carlini et al., 2024), we release a hash (sha512) with each mOSCAR image. mOSCAR is composed of 303M documents (200B tokens, 1.15B images) from 163 languages. Figure 2 shows the distribution of images and tokens per document and their joint distribution. As shown in Figure 2(a), the mean and median number of images per document is 2 and 3.80.
#docs. 10M 5M 1M 500K 200K 50K 10K 5K 1K
#langs. 10 15 36 46 56 75 119 133 154
@@
Figure 2: Distributions of numbers of tokens and images per document.
@@ 4.1.1 Text content
By contruction, mOSCAR is diverse in terms of number of languages, so we focus on the diversity of mOSCAR’s English documents and compare against mmc4 (Zhu et al., 2023), OBELICS (Laurençon et al., 2023) and the English subset of WIT (Srinivasan et al., 2021). We compute the Vendi score (Friedman and Dieng, 2023) on a set of SimCSE embeddings (Gao et al., 2021) with a RoBERTa encoder (Liu et al., 2019) to evaluate the content diversity. Since embedding-based diversity metrics target content diversity well but are less relevant for lexical diversity (Tevet and Berant, 2021), we measure lexical diversity via the distinct $n$-gram ratio (Li et al., 2016). An analysis of the topics (Grootendorst, 2022) found in multiple languages of mOSCAR where we show diverse topics across languages can be found in Appendix A.2.
@@ 4.1.2 Image diversity
Table 3: Average Vendi score ($\pm$ standard error) of images sampled from different datasets.
We compute the Vendi Score on random samples of images for different datasets, comparing the images from English mOSCAR documents with those from Conceptual Captions (Changpinyo et al., 2021), LAION-400M (Schuhmann et al., 2021) and WIT (Srinivasan et al., 2021). We represent each image by its SigLIP [^19] We use siglip-base-patch16-224. (Zhai et al., 2023) embedding and compute the Vendi score on batches of size 65,536 and a total of 1M images for each dataset. In Table 3, we notice that the set of images in mOSCAR documents are more diverse than images from WIT documents but less diverse than LAION-400M.
@@
Table 4: Average Vendi score ($\pm$ standard error) of images sampled from mOSCAR (English vs. any language).
@@
We also compare the diversity of images from English documents and of images sampled from documents of any language (English included). We use multilingual SigLIP (Chen et al., 2023a) trained on WebLI (Chen et al., 2023b) to compute image embeddings used to get the Vendi score. We again use a batch of size 65,536 and a total of 3M images, and we do not sample multiple images from a same document. For the multilingual setting, we randomly sample 50 languages and an equal number of images for each language to build the batch. As we did not do any image deduplication across languages, we could expect to have less diversity in the multilingual setting. However, Table 4 shows that the set of images is on average more diverse when sampled from all documents than from English-only documents. This means that the distribution of images is not exactly the same across languages, potentially due to cultural differences.
@@
Table 5: Results averaged over all languages. Multi. OF refers to multilingual Open Flamingo, mOSCAR + cap. refers to the model trained on text-image pairs and mOSCAR while cap. only refers to the model trained only on text-image pairs. Bold is best result.
xFlickR&CO XM3600 xGQA MaXM MaRVL XVNLI Multi30K CoMMuTE
#shots
Multi. OF (35M) 0 19.07 8.73 25.08 19.64 49.77 33.01 22.70 63.75
4 34.32 20.59 31.90 23.90 49.67 36.07 22.79 63.65
mOSCAR + cap. 8 36.77 22.15 33.9 24.41 49.72 37.16 23.21 63.00
16 37.63 22.24 35.71 25.38 49.73 35.36 23.48 62.77
Multi. OF (35M) 0 9.39 4.67 19.81 14.63 49.71 32.78 26.99 56.75
4 7.68 2.99 25.68 16.12 49.72 33.51 26.99 53.27
WIT + cap. 8 8.91 3.63 27.06 16.81 49.74 32.77 26.99 55.33
16 9.74 4.14 28.14 16.34 49.74 33.63 26.99 54.04
Table 6: Results averaged over all languages and comparison between a model trained on WIT and a checkpoint of multilingual Open Flamingo trained on 35M mOSCAR documents (full model was trained on 50M mOSCAR documents). Both models were trained on 35M documents from their respective training datasets and 70M text-image pairs for fair comparison. Multi. OF (35M) refers to multilingual Open Flamingo trained on 35M documents. Bold is best result.
5 Training a multilingual multimodal language model
We train a multilingual Flamingo-like model on mOSCAR that we call multilingual Open Flamingo. As adding captioning data to training data has been shown to improve zero-shot performance (McKinzie et al., 2024), we additionally train on LAION-400M, which we re-captioned using BLIP (Li et al., 2022), filtered with CLIP score (Radford et al., 2021) and translated using distilled NLLB-600M (Visheratin, 2023) following the proportion of languages found in mOSCAR. We use Gemma-2B (Gemma Team et al., 2024) as the underlying language model and we train the model on 50M mOSCAR documents and 100M randomly sampled image-text pairs. We also train a model on 300M image-text pairs, a model trained on 35M WIT (Srinivasan et al., 2021) documents and 70M text-image pairs and a model trained on 50M mOSCAR documents from the English subset and 100M English image-text pairs as comparison baselines. We additionally compare with OpenFlamingo-3B-MPT (Awadalla et al., 2023) as the translate-test baseline. The full list of languages for training and the implementation details can be found in Appendix A.6.
#shots xFlickR&CO XM3600 xGQA MaXM XVNLI
Multilingual OF 0 29.64 42.57 34.24 36.58 34.62
4 51.47 77.98 37.91 38.13 33.59
mOSCAR + cap. 8 56.75 77.64 39.44 38.52 38.75
16 59.89 78.18 40.09 35.80 36.60
English OF 0 32.70 43.75 34.71 36.19 35.82
4 51.39 75.33 37.48 37.96 34.88
English mOSCAR + English cap. 8 51.44 77.73 39.64 38.35 36.86
16 59.24 78.38 40.36 37.35 37.11
Table 7: Results on the English subsets of the test sets and comparison between multilingual Open Flamingo and an Open Flamingo trained on the English subset of mOSCAR and English text-image pairs (English OF). Both models were trained on 50M documents from their respective training datasets and 100M text-image pairs for fair comparison. Bold is best result.
5.1 Evaluation setup